# Zero-Shot Large Language Models for Long Clinical Text Summarization with Temporal Reasoning

### v2

@@ 4.2 Evaluation
On MIMIC, where the output is a summary, we report performance using three metrics that capture string overlap, semantics, and concepts. ROUGE-L  Lin (2004) measures string overlap, while BERTScore Zhang et al. (2019) assesses maximum token pairwise similarity. We use SapBERT as the backend for BERTScore due to its superior performance in biomedical entity representation Liu et al. (2020). We also compute the Concept F-score by mapping text to Concept Unique Identifiers (CUIs) using the Unified Medical Language System (UMLS)Lindberg et al. (1993). Specifically, both ground truth summaries and predicted text are converted into CUIs using QuickUMLS, a UMLS-based concept extractorSoldaini and Goharian (2016). We choose these metrics because of findings from prior work:  Croxford et al. (2024) reported that these metrics achieved higher correlation with physician judgement on clinical text evaluation; Myers et al. (2024) reported CUI F-score using QuickUMLS, identifying CUIs from seven semantic types, such as ‘Disease or Syndrome’ and ‘Sign or Symptom’. We adopt the same CUI F-score evaluation as in Myers et al. (2024).
On EHRShot, given that the diagnosis prediction tasks are essentially binary classification, we report macro-average accuracy and F-scores.
@@ 5 Results
Setting Model CUI ROUGE-L SapBERT Avg.
Direct Gen Mistral 13.46 ±6.41 16.28 ±12.83 65.23 ±13.38 31.66
Llama3 10.47 ±7.88 11.82 ±14.42 55.35 ±22.48 25.88
Qwen 15.13 ±6.44 15.28 ±5.38 63.98 ±13.50 31.46
DeepSeek 14.67 ±7.12 13.51 ±5.04 63.64 ±13.86 30.61
Llama2 15.79 ±7.38 15.34 ±5.90 63.82 ±11.44 31.65
RAG Mistral 12.15 ±2.39 15.04 ±2.34 65.89 ±5.69 31.03
Llama3 16.43 ±5.46 15.90 ±2.00 64.38 ±4.81 32.27
Qwen 16.90 ±4.36 17.91 ±2.35 65.25 ±4.77 33.35
Llama2 15.36 ±4.50 16.98 ±1.59 67.20 ±6.21 33.18
@@
Direct Gen RAG
Mistral Qwen Mistral Qwen
Acute MI Accuracy 52.84 62.72 63.03 65.50
F1 33.45 2.58 11.83 16.87
Celiac Disease Accuracy 62.96 96.05 95.04 95.30
F1 2.60 0.00 0.00 9.52
Hyperlipidemia Accuracy 58.02 68.15 66.09 63.03
F1 30.33 1.53 12.74 7.45
Hypertension Accuracy 64.44 69.38 68.81 68.81
F1 26.53 6.06 8.7 8.7
Lupus Accuracy 66.42 95.06 94.79 94.81
F1 10.53 0.00 0.00 0.00
Pancreatic Cancer Accuracy 73.83 78.52 78.86 79.05
F1 13.11 2.25 0.00 0.00
Discharge Summarization
Table 2 presents results from comparison between the Direct Generation and RAG approaches. Qwen achieves the highest average score of 33.35 using RAG. In general, most LLMs record a performance increase in their RAG variants, the biggest of which is Llama3’s, with an increase of 6.39 on average metric. Mistral is the exception, with a minor performance decrease of -0.63. DeepSeek, as the largest LLM, reports mediocre performance.
Table 3 presents the section-wise analysis of Llama2 and Mistral on direct generation, the two best LLMs directly handling long input. Diagnosis prediction, which demands the most clinical reasoning, shows the worst performance for both LLMs, highlighting the challenge of reasoning over patient trajectories.
Assessment and Plan Generation
Table 4 shows that adding prior context (Single-Day or Multi-Day) improves performance across all models, particularly for CUI and ROUGE scores. Llama3 consistently outperforms other models, especially in Single-Day and Multi-Day Contexts, reaching the highest Average scores across both Direct Generation and RAG settings. However, Multi-Day Context does not always yield performance gains over Single-Day, suggesting models struggle with longer patient histories. RAG generally improves performance, particularly for Mistral and Qwen, but its impact varies across models: Llama3 performs better without retrieval.
This task specifically requires clinical reasoning over past captured data (evidence) to summarize patient progression and plan for treatment, yet the relatively low ROUGE and CUI scores indicate that models still struggle with temporal reasoning and integrating historical context effectively.
Diagnosis Prediction
We evaluate Mistral and Qwen for this task based on their prior performance. Both models prioritize majority class predictions, leading to inflated accuracy but severely low F1-scores. RAG improves accuracy but does not meaningfully enhance recall or F1-score.
Overall Performance
6 Discussion
Another critical aspect to consider is running time and memory usage. All 7B LLMs require approximately 15GB of GPU RAM for direct generation. RAG demands significantly more memory, peaking at 32–33GB, nearly twice that of direct generation. For Qwen, it takes 18-20 minutes to run RAG on one progress note with prior context input, while directly generation takes 10 minutes on average. On EHRShot where input is shorter, diagnosis prediction takes just 3–4 seconds. For other LLMs with smaller context window, they run quicker.
Expert error analysis
Overall, RAG does slightly better, hallucinating less and generating more specific and clinically relevant summaries. However, both approaches still face significant issues: the models tend to focus on less relevant diagnoses and interpret them as the primary reason for admission and oversimplified summaries cause the omission of critical parts of the clinical course. The LLMs also struggle with diagnoses which are stated earlier on but are later discarded or disproved: they fail to realize that these diagnoses are no longer relevant or accurate. This emphasizes the models’ continuing struggle with temporal reasoning.
7 Conclusion
Our work examined LLM’s capabilities on long-context clinical summarization using two publicly available EHR datasets. We found that current LLMs continue to face significant challenges when it comes to accurately summarizing long-context clinical documents. Additionally, these models struggle with temporal reasoning over patient trajectories, making it difficult to track and interpret the sequence of medical events effectively. While RAG has shown some performance gains, the overall results remain substandard, emphasizing the need for further advancements in this area.
Limitations
Ethical Statement
Appendix A Appendix
A.1 More Preprocessing Details
A.2 EHRSHOT Statistics
A.3 Hyperparameter searching results
@@
A.4 Prompt and query optimization
@@
Table 9,10,12,13,11 presents all prompts and queries for MIMIC tasks.
A.5 Impact of context length on consecutive patient data
@@
A.6 More results for using ground-truth progress notes vs. generated progress notes on A&P Generation
@@ 1 Strongly Disagree; 2 Disagree; 3 Neutral; 4 Agree; 5 Strongly Agree
A.7 Break-down results behind Figure 4
@@
A.8 Details about human evaluation

### v3

@@ 4.2 Evaluation
On MIMIC, we report standard summarizatoin evaluation metrics that capture string overlap and semantics similarity. ROUGE-L  Lin (2004) measures string overlap, while BERTScore Zhang et al. (2019) assesses maximum token pairwise similarity. We use SapBERT as the backend for BERTScore due to its superior performance in biomedical entity representation Liu et al. (2020). On EHRShot, given that the diagnosis prediction tasks are essentially binary classification, we report macro-average accuracy and F-scores.
@@ 5 Results
Setting Model ROUGE-L BERTScore
Direct
Gen
Mistral 16.28 ±12.83 65.23 ±13.38
Llama3 11.82 ±14.42 55.35 ±22.48
Qwen 15.28 ±5.38 63.98 ±13.50
DeepSeek 13.51 ±5.04 63.64 ±13.86
Llama2 15.34 ±5.90 63.82 ±11.44
RAG Mistral 15.04 ±2.34 65.89 ±5.69
Llama3 15.90 ±2.00 64.38 ±4.81
Qwen 17.91 ±2.35 65.25 ±4.77
Llama2 16.98 ±1.59 67.20 ±6.21
@@
Section Method ROUGE-L BERTScore
Dx Direct Gen 3.42 ± 4.34 50.07 ± 10.78
CoT Prompt 2.95 ± 3.48 48.46 ± 11.00
HC Direct Gen 12.28 ± 3.13 62.51 ± 5.95
CoT Prompt 9.98 ± 3.84 59.60 ± 7.32
DI Direct Gen 12.07 ± 3.90 60.05 ± 9.09
CoT Prompt 10.83 ± 4.58 60.52 ± 7.86
Table 3: Mistral performance on the three sections of discharge summarization (Dx: Diagnosis, HC: Hospital course, DI: Discharge Instruction), comparing direct generation approach with the event extraction CoT.
Setting Direct Generation RAG
Mistral Llama3 Qwen Mistral Llama3 Qwen
RL BS RL BS RL BS RL BS RL BS RL BS
[…]
Table 4: Performance of Direct Generation and RAG methods on A&P generation across models and input settings. RL = ROUGE-L score; BS = BERTScore computed using SapBERT embeddings.
Model Metric Current Day Forward+N Backward-1
[…]
Table 5: Event extraction CoT results on A&P generation task across settings.
Direct Gen RAG
Mistral Qwen Mistral Qwen
Acute MI Accuracy 52.84 62.72 63.03 65.50
F1 33.45 2.58 11.83 16.87
Celiac Disease Accuracy 62.96 96.05 95.04 95.30
F1 2.60 0.00 0.00 9.52
Hyperlipidemia Accuracy 58.02 68.15 66.09 63.03
F1 30.33 1.53 12.74 7.45
Hypertension Accuracy 64.44 69.38 68.81 68.81
F1 26.53 6.06 8.7 8.7
Lupus Accuracy 66.42 95.06 94.79 94.81
F1 10.53 0.00 0.00 0.00
Pancreatic Cancer Accuracy 73.83 78.52 78.86 79.05
F1 13.11 2.25 0.00 0.00
Discharge Summarization. Table 2 presents results from comparison between the Direct Generation and RAG approaches. Qwen achieves the highest ROUGE-L score of 17.91 using RAG. In general, most LLMs record a performance increase in their RAG variants, the biggest of which is Llama3’s, with an increase of 4.08 on ROUGE-L. Mistral is the exception, with a minor performance decrease of -1.24. DeepSeek, as the largest LLM, reports mediocre performance
Table 3 shows section-wise analysis for Mistral comparing direct generation and CoT methods, where CoT event extraction performs slightly worse than direct generation across most sections. Other LLMs exhibit similar trends, so we report Mistral only for brevity. In Appendix 7, we provide more section-wise analysis on the Discharge Summ task. Results further confirm the challenge of LLMs reasoning over patient trajectories.
A& P Generation. Table 4 shows that adding prior context (Single- or Multi-Day) improves performance across all models, particularly for ROUGE scores. Llama3 performs best on direct generation, especially in Single- and Multi-Day Contexts, but is outperformed by Qwen and Mistral on RAG. However, Multi-Day Context does not always yield performance gains over Single-Day, suggesting models struggle with longer patient histories. RAG generally improves performance, particularly for Mistral and Qwen, but its impact varies across models: Llama3 performs better without retrieval.
This task specifically requires clinical reasoning over past captured data (evidence) to summarize patient progression and plan for treatment, yet the relatively low ROUGE and BERTScore values indicate that models still struggle with temporal reasoning and integrating historical context effectively.
Table 5 presents results of the event extraction CoT method. CoT did not yield improvements over direct generation or RAG approaches across ROUGE-L or BERTScore. This suggests that, while the event extraction step offers interpretability and control, it may not directly enhance generation quality when applied in a pipeline without further model adaptation. We view this as a useful diagnostic approach that could be further refined or integrated with instruction tuning in future work.
Diagnosis Prediction. We evaluate Mistral and Qwen for this task based on their prior performance. Both models prioritize majority class predictions, leading to inflated accuracy but severely low F1-scores. RAG improves accuracy but does not meaningfully enhance recall or F1-score.
Overall Performance
6 Discussion
Another critical aspect to consider is running time and memory usage. All 7B LLMs require approximately 15GB of GPU RAM for direct generation. RAG demands significantly more memory, peaking at 32–33GB, nearly twice that of direct generation. For Qwen, it takes 18-20 minutes to run RAG on one progress note with prior context input, while direct generation takes 10 minutes on average. On EHRShot where input is shorter, diagnosis prediction takes just 3–4 seconds. For other LLMs with smaller context window, they run quicker.
Expert error analysis
Overall, RAG performs slightly better, with fewer hallucinations and more clinically relevant summaries. However, both methods still struggle—often prioritizing less relevant diagnoses, oversimplifying summaries, and failing to discard outdated or disproven diagnoses. These issues highlight ongoing challenges with temporal reasoning.
7 Conclusion
We evaluated LLMs on long-context clinical summarization and prediction using two public EHR datasets. Current models struggle with accurate summarization and temporal reasoning, making it hard to interpret medical event sequences. While RAG offers some improvement, overall performance remains inadequate, highlighting the need for further progress.
Limitations
Ethical Statement
Appendix A Appendix
A.1 More Preprocessing Details
A.2 EHRShot Statistics
@@
A.3 More results on direct generation, RAG and CoT performance comparison
@@
Section Method CUI ROUGE-L BERTScore
Dx Direct Gen 5.46 ± 10.07 3.42 ± 4.34 50.07 ± 10.78
CoT Prompt 2.80 ± 5.04 2.95 ± 3.48 48.46 ± 11.00
HC Direct Gen 8.64 ± 5.98 12.28 ± 3.13 62.51 ± 5.95
CoT Prompt 7.78 ± 5.89 9.98 ± 3.84 59.60 ± 7.32
DI Direct Gen 9.09 ± 6.90 12.07 ± 3.90 60.05 ± 9.09
CoT Prompt 9.42 ± 8.19 10.83 ± 4.58 60.52 ± 7.86
Table 9: Mistral performance on the three sections of discharge summarization (Dx: Diagnosis, HC: Hospital course, DI: Discharge Instruction), comparing direct generation approach with the event extraction CoT. Including CUI f-score.
A.4 Hyperparameter searching results
@@
A.5 Prompt and query optimization
A.6 Event extraction COT prompt for discharge summarization
Chain-of-Thought Prompt for 48h Discharge Summary Event Extraction DISCHARGE EVENT EXTRACTION TASK Analyze the following data from the final 48 hours of the hospital stay and identify key clinical events that are most relevant for summarizing the course of treatment and informing discharge planning. {chronology_text} Only include events that reflect: 1. Significant changes in symptoms or status (e.g., improvements, worsening, new findings). 2. Clinically important test results (especially abnormal values that lead to certain treatments). 3. Major treatments or interventions (e.g., medication changes, procedures, escalation/de-escalation). 4. Care team decisions that indicate readiness for discharge or change in care goals. 5. Events linked to the final diagnosis or that inform follow-up care. Response Format: ### Day X Key Events ### - [Time]: [Description] (Reasoning)
Example Model Output ### Day 1 Key Events ### - **2159-03-12 08:00**: Stable Vital Signs (BP 120/70 mmHg, HR 88 bpm, $R$ 18 breaths/min, Temp 98.6°F). This indicates overall stability and readiness for discharge.
Figure 9: Chain-of-Thought prompting template and example output used for discharge summary event extraction based on the final 48 hours of hospitalization.
A.7 RAG and direct generation prompt optimization
@@
Table 13,14,16,17,15 presents all prompts and queries for MIMIC tasks.
A.8 Impact of context length on consecutive patient data
@@
A.9 More results for using ground-truth progress notes vs. generated progress notes on A&P Generation
@@ 1 Strongly Disagree; 2 Disagree; 3 Neutral; 4 Agree; 5 Strongly Agree
A.10 Break-down results behind Figure 5
@@
A.11 Details about human evaluation